# FSGS: Real-Time Few-shot View Synthesis using Gaussian Splatting

### 01 Dec 2023 v1

@@ Abstract
Novel view synthesis from limited observations remains an important and persistent task. However, high efficiency in existing NeRF-based few-shot view synthesis is often compromised to obtain an accurate 3D representation. To address this challenge, we propose a Few-Shot view synthesis framework based on 3D Gaussian Splatting that enables real-time and photo-realistic view synthesis with as few as three training views. The proposed method, dubbed FSGS, handles the extremely sparse initialized SfM points with a thoughtfully designed Gaussian Unpooling process. Our method iteratively distributes new Gaussians around the most representative locations, subsequently infilling local details in vacant areas. We also integrate a large-scale pre-trained monocular depth estimator within the Gaussians optimization process, leveraging online augmented views to guide the geometric optimization towards an optimal solution. Starting from sparse points observed from limited input viewpoints, our FSGS can accurately grow into unseen regions, comprehensively covering the scene and boosting the rendering quality of novel views. Overall, FSGS achieves state-of-the-art performance in both accuracy and rendering efficiency across diverse datasets, including LLFF, Mip-NeRF360, and Blender. Project website: https://zehaozhu.github.io/FSGS/.
@@ 1 Introduction
Novel view synthesis (NVS) from a set of view collections, as demonstrated by recent works [46, 35, 16], has played a critical role in the domain of 3D vision and is pivotal in many applications, e.g., VR/AR and autonomous driving. Despite its effectiveness in photo-realistic rendering, the requirement of dense support views has hindered its practical usages [33]. Previous studies have focused on reducing the view requirements by leveraging Neural Radiance Field (NeRF)[30], a powerful implicit 3D representation that captures scene details, combined with volume rendering techniques[13]. Depth regularizations [33, 54, 11, 47] within the density field, additional supervision from 2D pre-trained models [50, 23], large-scale pre-training [8, 59], and frequency annealings [55] have been proposed and adopted to address the challenge of few-shot view synthesis. However, NeRF-based methods incur high training and rendering costs to achieve satisfactory rendering quality. Though some NeRF follow-ups have reduced the training time in real-world scenes from days to hours [43, 51, 40, 58], or even minutes [31], a significant gap remains between achieving real-time rendering speed and photo-realistic high-resolution rendering quality.
@@
Figure 2: FSGS Pipeline. 3D Gaussians are initialized from SfM, with a few images (black cameras). For the sparsely placed Gaussians, we propose densifying new Gaussians to enhance scene coverage by unpooling existing Gaussians into new ones, with properly initialized Gaussian attributes. Monocular depth priors, enhanced by sampling unobserved views (red cameras), guide the optimization of grown Gaussians towards a reasonable geometry. The final loss consists of a photometric loss term, and a geometric regularization term calculated as depth correlation.
3D Gaussian Splatting (3D-GS)[26] has recently emerged as an efficient representation to model the 3D scenes from a dense collection of camera viewpoints. It represents the scene as a combination of 3D Gaussians with attributes for modeling intricate shape and appearances, and 2D images are rendered via splatting-based rasterization[57]. By replacing volume rendering with the efficient differentiable splatting, 3D Gaussian Splatting attains real-time rendering speed while maintaining the ability to render photo-realistic images from any new viewpoint. However, 3D-GS is initialized from Structure-from-Motion (SfM) points, and its performance strongly relies on both the quantity and accuracy of the initialized points. Although the subsequent Gaussian densification [26] can increase the number of Gaussians in both under-reconstructed and over-reconstructed regions, this straightforward strategy falls short in few-shot settings: it suffers from inadequate initialization, leading to oversmoothed outcomes and a tendency to overfit on training views.
Our goal is to construct a compact and efficient representation while preserving the visual fidelity from limited observations. To this end, we present FSGS, to model large-scale scenes from Sparse viewpoint inputs using 3D Gaussians as representation. To address the challenge of mitigating the gap in dense coverage from sparse input, the first challenge is how to move the locations of Gaussians effectively to cover and represent the 3D scene. Therefore, we propose Proximity-guided Gaussian Unpooling, where it grows new Gaussians by measuring the proximity of existing Gaussians with their neighbors. By placing new Gaussians in representative locations and initializing them with the information from existing Gaussians, it effectively increases the Gaussian count for comprehensive scene coverage. The next challenge is how to ensure the grown Gaussians correctly represent the geometry of scenes, even when there are not enough multi-view cues? Thus it becomes essential to leverage extra priors to control and regularize the optimization of the new Gaussians. Specifically, we propose to utilize rich depth priors with pre-trained monocular depth estimators on both training and online augmented pseudo views, which guide the Gaussian Unpooling to converge to a reasonable solution and ensure the geometric smoothness of the scenes. The back-propagation of the integrated depth prior is achieved by implementing a differentiable depth rasterizer. Validations on few-shot NVS benchmarks: scene-level LLFF, Mip-NeRF360, and object-level Blender datasets, demonstrates that FSGS achieves the state-of-the-art rendering quality, while can run at 203 FPS, a practical speed for real-world applications. Our contributions are outlined below:
@@
  • We propose a novel point-based framework for few-shot view synthesis, featuring Proximity-guided Gaussian Unpooling to densify Gaussians for comprehensive scene coverage.
  • Our framework integrates monocular depth priors, enhanced by virtually sampled training views, to guide the Gaussian optimization toward an optimal solution.
  • Equipped with an enhanced training paradigm, FSGS not only achieves real-time rendering speed (200+FPS) but also improves visual quality, paving the way for practical usage in real-world scenarios.
@@ 2.1 Neural Representations for 3D Reconstruction
The recent advancement of neural rendering techniques, such as Neural Radiance Fields (NeRFs) [30], has shown encouraging progress for novel view synthesis. NeRF learns an implicit neural scene representation that utilizes a MLP to map 3D coordinates $(x,y,z)$ and view dependency $(\theta,\phi)$ to color and density through a volume rendering function. Tremendous works focus on improving its efficiency [15, 31, 38, 43, 26, 7, 17], quality [48, 1, 3, 20, 42, 9, 52, 4], generalizing to unseen scenes [59, 8, 53, 24, 44, 10], applying artistic effects [14, 49, 22, 60] and 3D generation [34, 27, 28, 25, 45, 41, 21, 5, 18, 6]. In particular, Reiser et al. [38] accelerate NeRF’s training by splitting a big MLP into thousands of tiny MLPs. MVSNeRF [8] constructs a 3D cost volume [56, 19] and renders high-quality images from novel viewpoints. Moreover, Mip-NeRF [1] adopts conical frustum rather than a single ray in order to mitigate aliasing. Mip-NeRF 360 [2] further extends it to the unbounded scenes. While these NeRF-like models present strong performance on various benchmarks, they generally require several hours of training time. Muller et al. [31] adopt a multiresolution hash encoding technique that reduces the training time to 5 seconds. Kerbl et al. [26] propose to use a 3D Gaussian Splatting pipeline that achieves real-time rendering for either objects or unbounded scenes. The proposed FSGS approach is based on the 3D Gaussian Splatting framework but largely reduces the required training views.
@@ 2.2 Novel-View Synthesis Using Sparse Views
The original neural radiance field takes more than one hundred images as training views, which largely prohibits its practical usage. To tackle this issue, several works have attempted to reduce the number of training views. Specifically, Deng et al. [12] applies additional depth supervision to improve the rendering quality. Niemeyer et al. [32] proposes a depth smoothness loss as geometry regularization to stabilize training. DietNeRF [23] adds supervision on high-dimensional semantic space, the CLIP embedding space [36], to constraint the rendered unseen views. PixelNeRF [59] trains a convolution encoder to capture context information and learns to predict 3D representation from a single input. More recently, FreeNeRF [55] proposes a dynamic frequency controlling module that successfully trains a NeRF using sparse views with minimal modification. SparseNeRF [50] proposes a new spatial continuity loss to distill spatial coherence from monocular depth estimators. Compared to the aforementioned approaches, our method achieves real time rendering in novel views, and significantly improves the rendering quality.
@@ 3 Method
An overview of the FSGS framework is provided in Fig. 2. FSGS inputs with a limited number of images which are captured within a static scene. The camera poses and sparse point clouds are calculated from Structure-from-Motion(SfM) [39]. The initial 3D Gaussians used for further training are initialized from SfM points, with attributes of color, position and shape. The challenge in extremely sparse SfM points and insufficient observations is addressed by adopting Proximity-guided Gaussian Unpooling to densify Gaussians and fill the empty space by measuring the proximity between existing Gaussians and strategically placing new ones to the most representative location, to increase the capacity of handling details. To guarantee that the densified Gaussians can be optimized toward the correct scene geometry, we leverage the priors from 2D monocular depth estimator, enhanced by pseudo view generation which prevents our model from overfitting to sparse input viewpoints.
@@ 3.1 Preliminary and Problem Formulation
To ensure that $\Sigma$ is positive semi-definite and holds practical physical significance, $\Sigma$ is decomposed into two learnable components by $\Sigma={R}{S}{S}^{T}{R}^{T}$, where $R$ is a quaternion matrix representing rotation and $S$ is a scaling matrix
Figure 3: Sparse SfM Points vs. Synthesized Quality. The SfM points from COLMAP using 3-views (Bottom Left) is significantly sparse than full-view(Top Left). 3D-GS with sparse SfM points will decrease its quality when the training view number decreases.
In addition, each Gaussians store an opacity logit $o\in\mathbb{R}$ and the appearance feature represented by $n$ spherical harmonic (SH) coefficients $\left\{c_{i}\in{}^{3}|i=1,2,...,n\right\}$ where $n=D^{2}$ is the number of coefficients of SH with degree $D$. To render the 2D image, 3D-GS orders all the Gaussians that contributes to a pixel and blends the ordered Gaussians overlapping the pixels using the following function:
@@
where $c_{i}$ is the color computed from the SH coefficients of the $i^{\rm{th}}$ Gaussian. $\alpha_{i}$ is given by evaluating a 2D Gaussian with covariance $\Sigma^{\prime}\in{}^{2\times 2}$ multiplied by the opacity. The 2D covariance matrix $\Sigma^{\prime}$ is calculated by $\Sigma^{\prime}=JW\Sigma{W}^{T}{J}^{T}$, projecting the 3D covariance $\Sigma$ to the camera coordinates. Here, $J$ denotes the Jacobian of the affine approximation of the projective transformation, $W$ is the view transformation matrix.
A heuristic Gaussian densification scheme is introduced in 3D-GS [26], where Gaussians are densified based on an average magnitude of view-space position gradients which exceed a threshold. Although this method is effective when initialized with comprehensive SfM points, it is insufficient for fully covering the entire scene with an extremely sparse point cloud, from sparse-view input images. Additionally, some Gaussians tend to grow towards extremely large volumes, leading to results that overfit the training views and generalize badly to novel viewpoints (See Fig. 3).
@@ 3.2 Proximity-guided Gaussian Unpooling
Proximity Score and Graph Construction
During Gaussian optimization, we construct a directed graph, referred to as the proximity graph, to connect each existing Gaussian with its nearest $K$ neighbors by computing the Euclidean distance. We denote the originating Gaussian at the head as the “source” Gaussian, while the one at the tail as the “destination” Gaussian, which is one of the source’s $K$ neighbors. The proximity score assigned to each Gaussian is calculated as the average distance to its $K$ nearest neighbors. The proximity graph is updated following the densification or pruning process during optimization. We set $K$ to 3 in practice.
Gaussian Unpooling
@@
Figure 4: Gaussian Unpooling Illustration. We show a 2D toy case for visualizing Gaussian Unpooling with depth guidance, where the example 1D depth can provide priors on the relative distance of the Gaussians from the viewing direction, guide the Gaussian deformation toward a better solution.
@@ 3.3 Geometry Guidance for Gaussian Optimization
Having achieved dense coverage by unpooling Gaussians, a photometric loss with multi-view clues is applied for optimizing Gaussians. However, the insufficient observations in the sparse-view setting limit the capacity for learning coherent geometry, leading to a high risk of overfitting on training views and poor generalization to novel views. This necessitates the incorporation of additional regularization and priors to guide the Gaussian optimization. Specifically, we seek help from depth priors produced by a well-trained monocular depth estimator to guide the geometry of Gaussians towards a reasonable solution.
Injecting Geometry Coherence from Monocular Depth
We generate the monocular $\boldsymbol{D}_{est}$ depth maps at training views by using the pre-trained Dense Prediction Transformer (DPT), trained with 1.4 million image-depth pairs as a handy yet effective choice. To mitigate the scale ambiguity between the true scene scale and the estimated depth, we introduce a relaxed relative loss, Pearson correlation, on the estimated and rendered depth maps. It measures the distribution difference between 2D depth maps and follows the below function:
Differentiable Depth Rasterization
To enable the backpropogation from depth prior to guide Gaussian training, we implement a differentiable depth rasterizer, allowing for receiving the error signal between the rendered depth $\boldsymbol{D}_{ras}$ and the estimated depth $\boldsymbol{D}_{est}$. Specifically, we utilize the alpha-blending rendering in 3D-GS for depth rasterization, where the z-buffer from the ordered Gaussians contributing to a pixel is accumulated for producing the depth value:
$d=\sum_{i=1}^{n}d_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (4)
Here $d_{i}$ represents the z-buffer of the $i^{\rm{th}}$ Gaussians and $\alpha$ is identical to that in Eq. 2. The fully differentiable implementation enables the depth correlation loss, further improving the similarity between the rendered and estimated depths.
Synthesize Pseudo Views

### 16 Jun 2024 v2

@@ Abstract
Novel view synthesis from limited observations remains a crucial and ongoing challenge. In the realm of NeRF-based few-shot view synthesis, there is often a trade-off between the accuracy of the synthesized view and the efficiency of the 3D representation. To tackle this dilemma, we introduce a Few-Shot view synthesis framework based on 3D Gaussian Splatting, which facilitates real-time, photo-realistic synthesis from a minimal number of training views. FSGS employs an innovative Proximity-guided Gaussian Unpooling, specifically designed for sparse-view settings, to bridge the gap presented by the extremely sparse initial point sets. This method involves the strategic placement of new Gaussians between existing ones, guided by a Gaussian proximity score, enhancing the adaptive density control. We have identified that Gaussian optimization can sometimes result in overly smooth textures and a propensity for overfitting when training views are limited. To mitigate these issues, FSGS introduces the synthesis of virtual views to replicate the parallax effect experienced during training, coupled with geometric regularization applied across both actual training and synthesized viewpoints. This strategy ensures that new Gaussians are placed in the most representative locations, fostering more accurate and detailed scene reconstruction. Our comprehensive evaluation across various datasets—including NeRF-Synthetic, LLFF, Shiny, and Mip-NeRF360 datasets—illustrates that FSGS not only delivers exceptional rendering quality but also achieves an inference speed more than 2000 times faster than existing state-of-the-art methods for sparse-view synthesis. Project webpage https://zehaozhu.github.io/FSGS/.
Keywords:
Neural Rendering Gaussian Splatting Sparse View Synthesis
@@ 1 Introduction
Novel view synthesis (NVS) from a set of view collections, as demonstrated by recent works [46, 35, 16], has played a critical role in the domain of 3D vision and is pivotal in many applications, e.g., VR/AR and autonomous driving. Despite its effectiveness in photo-realistic rendering, the requirement of dense support views has hindered its practical usages [33]. Previous studies have focused on reducing the view requirements by leveraging Neural Radiance Field (NeRF)[30], a powerful implicit 3D representation that captures scene details, combined with volume rendering techniques[13]. Depth regularization [33, 56, 11, 47] within the density field, additional supervision from 2D pre-trained models [50, 23], large-scale pre-training [8, 61], and frequency annealings [57] have been proposed and adopted to address the challenge of few-shot view synthesis. While these NeRF-based approaches are promising, they often lead to substantial computational demands, which can affect real-time performance adversely. Subsequent research has managed to reduce training time in real-world scenarios from days to mere hours [43, 51, 40, 60], and even minutes in some cases [31]. However, a noticeable gap persists between attaining real-time rendering speeds and the desired photo-realistic, high-resolution output quality.
@@
In our research, we delve into the advancements in efficient 3D Gaussian Splatting (3D-GS) [26] and examine the challenges associated with deploying 3D-GS on sparse inputs. A crucial aspect for the effectiveness of 3D-GS is the densification process, which transforms the sparse initial point cloud into a more detailed representation of the 3D environment. However, the placement of new Gaussians, dictated by the spatial gradient, tends to be noisy and unrepresentative, especially in sparse-view scenarios. Additionally, the reliance on photometric loss with limited view counts often results in overly smooth textures when adhering to the conventional densification approach.
To address these issues, we introduce Proximity-guided Gaussian Unpooling, a novel strategy designed specifically for sparse inputs. This method enhances the Gaussian representation by inserting new Gaussians between existing ones, based on the proximity to their neighbors. This strategic placement, combined with the initialization using observations from existing Gaussians, significantly improves scene representation by increasing the Gaussian density. Furthermore, we advocate for view augmentation through the generation of virtual camera, not present during training, to apply additional constraints in sparse setups. Incorporating monocular depth priors helps regularize the new Gaussians, steering them towards a plausible solution while enhancing texture detail, essential for accurate relative positioning in both actual training and synthetic camera views. We have conducted thorough evaluations of our Few-Shot Gaussian Splatting (FSGS) framework across a variety of few-shot Novel View Synthesis benchmarks. These include the object-centric NeRF-Synthetic datasets, the forward-facing LLFF datasets, the Shiny datasets with intricate lighting conditions, and the unbounded Mip-NeRF360 datasets. Our experiments demonstrate that FSGS sets a new benchmark in rendering quality and operates at a real-time speed (203 FPS), making it suitable for real-world applications. The efficacy of our method allows FSGS to outperform 3D-GS even with fewer Gaussians, enhancing both the efficiency and quality on the rendered scenes.
Our key contributions are as follows:
@@
  • We propose a novel point-based framework, FSGS, for few-shot view synthesis that densifies new Gaussians via Proximity-guided Gaussian Unpooling. This method effectively increases the density of Gaussians, ensuring detailed and comprehensive scene representation.
  • FSGS addresses the overfitting challenge inherent in sparse-view Gaussian splatting. It achieves this by generating unseen viewpoints during training and incorporating distance correspondences on both training and synthesized pseudo views. This strategy directs the Gaussian optimization process toward solutions that are both highly accurate and visually compelling.
  • FSGS significantly enhances the visual quality, and also facilitates real-time rendering speeds (over 200 FPS) leading to a viable option for practical implementation in various real-world applications.
@@ 2.1 Neural Representations for 3D Reconstruction
The recent advancement of neural rendering techniques, such as Neural Radiance Fields (NeRFs) [30], has shown encouraging progress for novel view synthesis. NeRF learns an implicit neural scene representation that utilizes a MLP to map 3D coordinates $(x,y,z)$ and view dependency $(\theta,\phi)$ to color and density through a volume rendering function. Tremendous works focus on improving its efficiency [15, 31, 38, 43, 26, 7, 17], quality [48, 1, 2, 20, 42, 9, 52, 4], generalizing to unseen scenes [61, 8, 53, 24, 44, 10], applying artistic effects [14, 49, 22, 62] and 3D generation [34, 27, 28, 25, 45, 41, 21, 5, 18, 6]. In particular, Reiser et al. [38] accelerate NeRF’s training by splitting a big MLP into thousands of tiny MLPs. MVSNeRF [8] constructs a 3D cost volume [59, 19] and renders high-quality images from novel viewpoints. Moreover, Mip-NeRF [1] adopts conical frustum rather than a single ray in order to mitigate aliasing. Mip-NeRF 360 [3] further extends it to the unbounded scenes. While these NeRF-like models present strong performance on various benchmarks, they generally require several hours of training time. Muller et al. [31] adopt a multiresolution hash encoding technique that reduces the training time significantly. Kerbl et al. [26] propose to use a 3D Gaussian Splatting pipeline that achieves real-time rendering for either objects or unbounded scenes. The proposed FSGS approach is based on the 3D Gaussian Splatting framework but largely reduces the required training views.
@@ 2.2 Novel-View Synthesis Using Sparse Views
The original neural radiance field takes more than one hundred images as input, largely prohibiting its practical usage. To tackle this issue, several works have attempted to reduce the number of training views. Specifically, Depth-NeRF [12] applies additional depth supervision to improve the rendering quality. RegNeRF [32] proposes a depth smoothness loss as geometry regularization to stabilize training. DietNeRF [23] adds supervision on the CLIP embedding space [36], to constraint the rendered unseen views. PixelNeRF [61] trains a convolution encoder to capture context information and learns to predict 3D representation from sparse inputs. More recently, FreeNeRF [57] proposes a dynamic frequency controlling module for few-shot NeRF. SparseNeRF [50] proposes a new spatial continuity loss to distill spatial coherence from monocular depth estimators. Concurrent work ReconFusion [55] employs diffusion models to synthesize additional views, which may not always adhere to view consistency and time consuming. ReconFusion jointly train a Zip-NeRF with synthesized views under a sparse-view setting. In contrast, our method improves the optimization process of Gaussian Splatting, and facilitates both the real-time rendering speed and rendering quality.
@@ 3 Method
Overview.
This section provides an overview of the FSGS framework, as illustrated in Fig. 2. FSGS processes a limited set of images captured from a static scene. The camera poses and point clouds are derived using the Structure-from-Motion (SfM) software, COLMAP[39]. The initialization of 3D Gaussians is based on a sparse point cloud, incorporating attributes such as color, position, and a predefined conversion rule for shape and opacity. The issue of extremely sparse points is tackled through the implementation of Proximity-guided Gaussian Unpooling. This method densifies Gaussians and populates the empty spaces by assessing the proximity between existing Gaussians and positioning new ones in the most representative areas, thereby enhancing scene details. To mitigate overfitting in standard 3D-GS with sparse-view data, we introduce the generation of pseudo camera viewpoints around the training cameras. This approach, coupled with the geometry regularization, steers the model towards accurately reconstructing the scene’s geometry.
Figure 2: FSGS Pipeline. 3D Gaussians are initialized from COLMAP, with a few images (black cameras). For the sparsely placed Gaussians, we propose densifying new Gaussians to enhance scene coverage by unpooling existing Gaussians into new ones, with properly initialized Gaussian attributes. Monocular depth priors, enhanced by sampling unobserved views (red cameras), guide the optimization of grown Gaussians towards a reasonable geometry. The final loss consists of a photometric loss term, and a geometric regularization term calculated as depth relative correspondence.
@@ 3.1 Preliminary and Problem Formulation
To ensure that $\Sigma$ is positive semi-definite and holds practical physical significance, $\Sigma$ is decomposed into two learnable components by $\Sigma={R}{S}{S}^{T}{R}^{T}$, where $R$ is a quaternion matrix representing rotation and $S$ is a scaling matrix.
Each Gaussians also store an opacity logit $o\in\mathbb{R}$ and the appearance feature represented by $n$ spherical harmonic (SH) coefficients $\left\{c_{i}\in^{3}|i=1,2,...,n\right\}$ where $n=D^{2}$ is the number of coefficients of SH with degree $D$. To render the 2D image, 3D-GS orders all the Gaussians that contributes to a pixel and blends the ordered Gaussians overlapping the pixels using the following function:
@@
where $c_{i}$ is the color computed from the SH coefficients of the $i^{\rm{th}}$ Gaussian. $\alpha_{i}$ is given by evaluating a 2D Gaussian with covariance $\Sigma^{\prime}\in^{2\times 2}$ multiplied by the opacity. The 2D covariance matrix $\Sigma^{\prime}$ is calculated by $\Sigma^{\prime}=JW\Sigma{W}^{T}{J}^{T}$, projecting the 3D covariance $\Sigma$ to the camera coordinates. Here, $J$ denotes the Jacobian of the affine approximation of the projective transformation, $W$ is the view transformation matrix. A heuristic Gaussian densification scheme is introduced in 3D-GS [26], where Gaussians are densified based on an average magnitude of view-space position gradients which exceed a threshold. Although this method is effective when initialized with comprehensive SfM points, it is insufficient for fully covering the entire scene with an extremely sparse point cloud, from sparse-view input images. Additionally, some Gaussians tend to grow towards extremely large volumes, leading to results that overfit the training views and generalize badly to novel viewpoints (See Fig. 3).
However, 3D-GS is initialized from SfM points, and its performance strongly relies on both the quantity and accuracy of the initialized points. Although the subsequent Gaussian densification [26] can increase the number of Gaussians in both under-reconstructed and over-reconstructed regions, this straightforward strategy falls short in few-shot settings: it suffers from inadequate initialization, leading to oversmoothed outcomes and a tendency to overfit on training views.
Figure 3: Points Sparsity vs. Synthesized Quality. The SfM points from COLMAP using 3-views (Bottom Left) is significantly sparse than full-view(Top Left). 3D-GS with sparse SfM points will decrease its quality when the training view number decreases.
@@ 3.2 Proximity-guided Gaussian Unpooling
Proximity Score and Graph Construction.
During Gaussian optimization, we construct a directed graph, referred to as the proximity graph, to connect each existing Gaussian with its nearest $K$ neighbors by computing the proximity (a.k.a. Euclidean distance). Specifically, we denote the originating Gaussian at the head as the “source” Gaussian, while the one at the tail as the “destination” Gaussian, which is one of the source’s $K$ neighbors. These “destination” Gaussians are determined via the rule:
$D_{i}^{K}$ $=K\text{-min}(d_{ij}),\quad\forall j\neq i$ (3)
Here, $d_{ij}$ is calculated via $d_{ij}=\|\mu_{i}-\mu_{j}\|$, representing the Euclidean distance among the centers of Gaussian $G_{i}$ and Gaussian $G_{j}$. The assigned proximity score $P_{i}$ to Gaussian $G_{i}$ is calculated as the average distance to its $K$ nearest neighbors:
$P_{i}$ $=\frac{1}{K}\sum_{j=1}^{K}D_{i}^{K}$ (4)
The proximity graph is updated following the densification or pruning process during optimization. We set $K$ to 3 in practice.
Figure 4: Gaussian Unpooling Illustration. We show a 2D toy case for visualizing Gaussian Unpooling with depth guidance, where the example 1D depth provides priors on the relative distance of the Gaussians from the viewing direction, guide the Gaussian deformation toward a better solution.
Gaussian Unpooling.
@@ 3.3 Geometry Guidance for Gaussian Optimization
Having achieved dense coverage by unpooling Gaussians, a photometric loss with sparse-view clues is applied for optimizing Gaussians. However, the insufficient parallax in the sparse-view setting limit the 3D Gaussians to be optimized toward a globally consistent direction where it tend to overfit on training views, and poor generalization to novel views. To inject more regularization to the optimization, we propose to create some virtual cameras that unseen in training, and apply the pixel wise geometric correspondences as additional regularization.
Synthesize Pseudo Views.
@@
Inject Geometry Coherence from Monocular Depth.
We generate the monocular $\boldsymbol{D}_{est}$ depth maps at both training and pseudo views by using the pre-trained Dense Prediction Transformer (DPT) [37], trained with 1.4 million image-depth pairs as a handy yet effective choice. To mitigate the scale ambiguity between the true scene scale and the estimated depth, we introduce a relaxed relative loss, Pearson correlation, on the estimated and rendered depth maps. It measures the distribution difference between 2D depth maps and follows the below function:
Differentiable Depth Rasterization.
To enable the backpropogation from depth prior to guide Gaussian training, we implement a differentiable depth rasterizor, allowing for receiving the error signal between the rendered depth $\boldsymbol{D}_{ras}$ and the estimated depth $\boldsymbol{D}_{est}$. Specifically, we utilize the alpha-blending rendering in 3D-GS for depth rasterization, where the z-buffer from the ordered Gaussians contributing to a pixel is accumulated for producing the depth value:
$d=\sum_{i=1}^{n}d_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (7)
Here $d_{i}$ represents the z-buffer of the $i^{\rm{th}}$ Gaussians and $\alpha$ is identical to that in Eq. 2. This differentiable implementation enables the depth correlation loss.